OpenStudy (abbycross167):

Solve: 5|2x + 1 | = 55 My Answer~ Divide both side by 5: |2x+1|=11 2x+1=11 and 2x+1= -11 -1 -1 -1 -1 2x = 12 2x = -10 X=5 x=-6 What my teacher said~ 2x = 12 and 2x = 10 yes but your x values are incorrect ~~Can someone please help me fix what I did wrong?

pooja195 (pooja195):

because you flipped them.

pooja195 (pooja195):

My Answer~ Divide both side by 5: |2x+1|=11 2x+1=11 and 2x+1= -11 -1 -1 -1 -1 2x = 12 2x = -10 X=5 ---------> <------- x=-6

OpenStudy (abbycross167):

oh okay, thank you very much!

pooja195 (pooja195):

=)